### v1

# Predicting from a Different Perspective in Re-ranking Model for Inductive Knowledge Graph Completion

Abstract
Rule-induction models have been shown great power in the inductive setting of knowledge graph completion. In this setting, the models are tested on a knowledge graph entirely composed of unseen entities. These models learn relation patterns as rules by utilizing subgraphs. The same input but different rules cause differences in the model’s predictions. In this paper, we focus on this behavior of the model. We propose a re-ranking-based model called ReDistLP (Re-ranking with a Distinct Model for Link Prediction). This model enhances the effectiveness of re-ranking by leveraging the difference in the predictions between the initial retriever and the re-ranker. ReDistLP outperforms the state-of-the-art methods in 2 out of 3 datasets.
1 Introduction
Knowledge graphs (KGs) store collections of facts in the form of triples. Each triple $(h,r,t)$ represents a relation $r$ between a head entity $h$ and a tail entity $t$. Various KGs have been proposed, including Freebase (Bollacker et al., 2008) for real-world information and GeneOntology (Consortium, 2004) for biological data. These KGs are used in many applications, such as question answering and recommendation systems. Since many KGs are inherently incomplete, it is essential to complete the missing information. Link prediction is a task to predict a missing link in a KG. This completion process improves the KG itself and the usefulness of applications using KGs.
Knowledge Graph Embedding (KGE) such as TransE (Bordes et al., 2013), RotatE (Sun et al., 2019), and ConvE (Dettmers et al., 2018), is the most dominant method for link prediction. Most KGE methods are designed for the transductive setting, where the test entities are a subset of the training entities. These models may not produce meaningful embeddings for unseen entities in the inductive setting. In the inductive setting, entities for the test are not present in the training, while relations are the same (Teru et al., 2020). This setting requires the model to have high inductive ability. Inductive ability is the capability to generalize and make accurate predictions for unseen entities based on the learned patterns and rules from the training data.
GraIL (Teru et al., 2020) and BERTRL (Zha et al., 2022) are designed for the inductive setting. They learn the probabilistic logical rules of relation patterns using subgraphs around an input triple. These models demonstrate a high inductive ability because such rules are entity-independent.
Our empirical findings reveal that BERTRL tends to predict different entities when distinct rules are provided as input. The model trained with the longer rules predicts entities within narrower regions on a KG (Section 4). Motivated by this observation, we propose ReDistLP (Re-ranking with a Distinct Model for Link Prediction), which leverages this phenomenon to improve accuracy. ReDistLP introduces a re-ranking pipeline. The re-ranking pipeline obtains a pool of candidate documents through an inexpensive approach and subsequently re-evaluates them using a more expensive model (Wang et al., 2011). This approach achieves both low computational cost and high accuracy. While existing research has mainly focused on improving the recall of the initial retriever or the performance of the re-ranker, this study focuses on the difference in the predictions between them. We show theoretical insights to improve the accuracy of the re-ranking pipeline. When selecting a model as the re-ranker, choosing one that predicts different entities compared to the initial retriever can improve accuracy. We train three variants of BERTRL, each incorporating rules of different lengths as additional inputs during training and inference. We construct ReDistLP using these variants of BERTRL. ReDistLP achieves the best performance in 2 out of 3 datasets.
The rest of this paper is organized as follows. In Section 2, we describe some related work. In Section 3, we show theoretical insights for maximizing the effect of re-ranking. Section 4 analyzes how the model makes different predictions based on distinct rules. In Section 5, we present our ReDistLP model and provide a detailed explanation of its components. In Section 6, we evaluate the performance of ReDistLP on three benchmark datasets. Section 7 concludes the paper.
2 Related Works and Backgrounds
2.1 Problem Definition
Formally, a knowledge graph $G$ is defined as a tuple $(\mathcal{E},\mathcal{R},\mathcal{T})$, where $\mathcal{E}$ is a set of entities, $\mathcal{R}$ is a set of relations and $\mathcal{T}$ is a set of triples $\{(h_{i},r_{i},t_{i})\}$. Here, $h_{i},t_{i}\in\mathcal{E}$ denotes the subject and the object, respectively, $r_{i}\in\mathcal{R}$ denotes the relation (Teru et al., 2020).
In the inductive setting of link prediction, we have a training graph $G_{{\rm train}}=(\mathcal{E}_{{\rm train}},\mathcal{R},\mathcal{T}_{{\rm train}})$ and a inference graph $G_{{\rm inf}}=(\mathcal{E}_{{\rm inf}},\mathcal{R},\mathcal{T}_{{\rm inf}})$. The entity sets of $G_{{\rm train}}$ and $G_{{\rm inf}}$ are mutually exclusive ($\mathcal{E}_{train}\cap\mathcal{E}_{inf}=\emptyset$).
Formally, most models designed for the link prediction task assign scores to each entity in response to an input query such as $(h,r)$, and the higher score signifies a greater likelihood of a relevant or accurate link.
2.2 Knowledge Graph Completion Models
Embedding-based Models
Embedding-based models such as TransE (Bordes et al., 2013), RotatE (Sun et al., 2019), and ConvE (Dettmers et al., 2018), learn the embedding of entities and relations in continuous spaces. These methods are trained to ensure that the operation of the relation on the subject results in the object in the vector space. KG Embedding-based models capture the structural properties of the KG within this embedding space. However, most embedding-based models are transductive; They do not produce appropriate embeddings for unseen entities during the training. Some methods, such as those proposed by (Wang et al., 2019; Hamaguchi et al., 2017), generate embeddings for unseen entities by aggregating the embeddings of neighboring entities learned during training. However, these methods cannot be applied in the inductive setting that handles a KG entirely composed of unseen entities. MorsE (Chen et al., 2022) and StATIK (Markowitz et al., 2022) are embedding-based models that can predict in the inductive setting. MorsE introduces meta-learning to transfer knowledge to unseen entities. StATIK uses a transformer and a message-passing neural network as the encoder, with TransE as the decoder.
Rule-induction Methods and Subgraph-induction Methods
Rule-induction methods learn probabilistic logical rules from a KG. This method inherently can solve the inductive setting since such rules are independent of entities. Here is an example of rules
$(x,{\rm president\_of},y)\land(z,{\rm capital\_of},y)$
$\to(x,{\rm work\_at},z).$ (1)
This rule denotes that if $x$ is the president of $y$, and $z$ is the capital of $y$, then $x$ works at $z$. Once such rules are learned, they can be used to generalize unseen entities. AMIE (Galárraga et al., 2013) and RuleN (Meilicke et al., 2018) mind these rules while reducing the search space. NeuralLP (Yang et al., 2017) and DRUM (Sadeghian et al., 2019) learn such rules in an end-to-end differentiable manner. GraIL (Teru et al., 2020) uses a graph neural network to learn these rules over the subgraphs around the input triple.
Language-based Models
KG-BERT (Yao et al., 2019), GenKGC (Xie et al., 2022), and BERTRL (Zha et al., 2022) are language-based models that use pre-trained language models to utilize the textual information. KG-BERT takes the triple as the textual sequences. It trains BERT to score the plausibility of the input triple. GenKGC takes the missing triple $(h,r,?)$ and generates a missing entity. While KG-BERT requires scoring all possible triples in inference, GenKGC can infer entities directly through generation. Thus, GenKGC offers faster inference times than KG-BERT. However, both KG-BERT and GenKGC have limited inductive ability since they do not utilize subgraph information. Their inductive capacity is solely attributed to the pre-trained language model. In contrast, BERTRL demonstrates a high inductive ability. Similar to KG-BERT, BERTRL is a model designed to identify triples. This model takes the paths connecting head and tail entities, along with the input triple. Its high inductive capability arises from both a pre-trained language model and a subgraph reasoning mechanism. Hence, BERTRL can also be classified as a subgraph-induction method. The formulation for BERTRL is as follows
$score(h,r,t)=\max_{\rho}p(y=1|h,r,t,\rho),$ (2)
where $\rho$ denotes a path connecting entities $h$ and $t$. The score of an input triple is the maximum score obtained over all possible paths.
2.3 Re-ranking Models
The re-ranking pipeline initially retrieves a collection of candidate documents with a less costly model and subsequently re-scores them using a more expensive model (Wang et al., 2011). Usually, the re-ranker model has higher accuracy but entails higher inference costs. This trade-off is crucial for the efficiency and effectiveness of the model. A higher number of candidate documents need to be retrieved to increase the efficacy, but this leads to a decrease in efficiency. Reducing the size of the candidate set can enhance efficiency, as it enables the re-ranker to minimize the time spent on re-ranking the documents; however, this results in a reduction in effectiveness (Chen et al., 2017; MacAvaney et al., 2022).
Various re-ranking models have been proposed to enhance either efficiency or effectiveness. GAR (MacAvaney et al., 2022) dynamically updates the initial pool to improve the recall of the initial retriever. This model repeatedly adds a document similar to the initial pool’s top-scoring ones. The monoBERT (Nogueira et al., 2019) and duoBERT (Nogueira et al., 2019) are BERT-based textual ranking models. The monoBERT scores documents based on binary classification. The duoBERT guesses which sentence is more relevant to the query. (Chen et al., 2017) proposed a cost-aware re-ranking model by introducing the feature costs. This model optimizes latency by selecting feature sets based on the importance and costs of features.
3 Maximize the Effectiveness of Re-ranking
We provide theoretical insights to maximize the effectiveness of re-ranking. Let $score_{m}$ represent the score function in model $m$, and $\bar{t}\in\mathcal{E}$ denote the correct answer. Now we consider an entity set where the score of each entity is greater than or equal to the score of $\bar{t}$:
$S_{m}(h,r)=\{t\in\mathcal{E}|score_{m}(h,r,t)\geq score_{m}(h,r,\bar{t})\}.$ (3)
The size of $S_{m}(h,r)$ is equivalent to the performance of model $m$ for query $(h,r)$ because $1/|S_{m}(h,r)|$ corresponds to the Reciprocal Rank (RR). Mean Reciprocal Rank (MRR) is widely used as an evaluation metric in research fields such as information retrieval systems and recommender systems.
Suppose that model $m_{1}$ is a initial retriever, and model $m_{2}$ is a re-ranker. Let us assume the accuracy of $m_{2}$ exceeds that of $m_{1}$ ( $|S_{m_{1}}|>|S_{m_{2}}|$). After re-ranking, the predicted entities are given by
$S^{\prime}_{m_{1},m_{2}}(h,r)=\{t\in S_{m_{1}}(h,r)|score_{m_{2}}(h,r,t)\geq score_{m_{2}}(h,r,\bar{t})\}.$ (4)
Given the above, we can infer
$|S^{\prime}_{m_{1},m_{2}}(h,r)|\leq|S_{m_{2}}(h,r)|<|S_{m_{1}}(h,r)|.$ (5)
This equation indicates that the performance after re-ranking exceeds that of the initial retriever and the re-ranker. Equation (5) can be resolved into the following equation and inequality, depending on the relationship between $S_{m_{1}}$ and $S_{m_{2}}$:
$S_{m_{1}}\subset S_{m_{2}}\implies|S^{\prime}_{m_{1},m_{2}}|=|S_{m_{2}}|.$ (6)
This equation suggests that when $S_{m_{1}}$ is a subset of $S_{m_{2}}$, there will be no improvement in accuracy after re-ranking:
$S_{m_{1}}\not\subset S_{m_{2}}\implies|S^{\prime}_{m_{1},m_{2}}|<|S_{m_{2}}|.$ (7)
This equation implies that when $S_{m_{1}}$ is not a subset of $S_{m_{2}}$, the accuracy after re-ranking will be improved compared to that of the re-ranker.
Furthermore, by selecting models based on the next equation, re-ranking effectiveness is maximized through the minimization of $S_{m_{1},m_{2}}$:
$\operatorname*{argmax}_{\;\;m_{1},m_{2}}\;|S_{m_{1}}\setminus S_{m_{2}}|.$ (8)
From the above observations, it is ineffective to select the models that make almost the same predictions with each other; it is crucial to choose the models that offer distinct predictions from one another.
However, in practical use, no guarantee that the correct answer $\bar{t}$ is in the initial pool. This is because the initial retriever selects from top-$k$ instances. Therefore, it is important to maximize the effectiveness of re-ranking, considering both the performance of the initial retriever $m_{1}$ and the relationship between models in practical use.
4 Analyzing Model Behavior Under Different Rule Sets
In this section, we demonstrate that BERTRL predicts distinct entities when provided with different rules as input. We train BERTRL with 3-hop, 2-hop, and 1-hop rules, respectively, and analyze the differences in their prediction sets.
To investigate the differences in predictions between the variant models directly, we summarize the intersection ratio of top-$n$ prediction sets in Table 1. The results show that the prediction sets differ significantly depending on the number of hops used in the training rules. The predictions of BERTRL trained with 1-hop and 2-hop rules are the most similar, while those trained with 1-hop and 3-hop rules are the least similar.
Table 1: Intersection Ratio of Top-$n$ Prediction Sets
WN18RR FB15k-237 NELL-995
$n=10$
1-hop 2-hop 0.0711 0.372 0.471
1-hop 3-hop 0.0364 0.274 0.354
2-hop 3-hop 0.0549 0.342 0.430
$n=100$
1-hop 2-hop 0.236 0.431 0.606
1-hop 3-hop 0.104 0.282 0.476
2-hop 3-hop 0.103 0.294 0.514
$n=1000$
1-hop 2-hop 0.921 0.970 0.790
1-hop 3-hop 0.921 0.970 0.790
2-hop 3-hop 0.921 0.970 0.790
Table 2: $n$-hop Reachability Between Entities in Top-10 Predicted Sets
WN18RR FB15k-237 NELL-995
$n=2$
BERTRL 3-hop 0.338 0.639 0.894
BERTRL 2-hop 0.178 0.612 0.899
BERTRL 1-hop 0.0624 0.283 0.538
$n=3$
BERTRL 3-hop 0.441 0.860 0.980
BERTRL 2-hop 0.239 0.777 0.958
BERTRL 1-hop 0.110 0.487 0.740
$n=5$
BERTRL 3-hop 0.577 0.958 0.996
BERTRL 2-hop 0.392 0.910 0.992
BERTRL 1-hop 0.282 0.797 0.959
Furthermore, we observe that the longer the number of hops, the more the predicted entities tend to be concentrated locally on the graph. Figure 1 shows the top-10 predicted entities and the paths that connect them within 3-hop. Table 2 shows $n$-hop reachability between entities in top-10 predicted sets.
(a) Hop 1 (WN18RR)
(b) Hop 1 (FB15k-237)
(c) Hop 1 (NELL-995)
(d) Hop 2 (WN18RR)
(e) Hop 2 (FB15k-237)
(f) Hop 1 (NELL-995)
(g) Hop 3 (WN18RR)
(h) Hop 2 (FB15k-237)
(i) Hop 3 (NELL-995)
Figure 1: Visualization of predicted top-10 entities. Red nodes represent entities from the input triple. Green nodes represent predicted entities. Blue nodes represent entities that appear in connected paths.
From Table 2 and Figure 1, it can be observed that: (1) Reachability in predicted entities increases with the number of hops of rules. (2) BERTRL with a higher number of hops predicts entities within narrower regions on a KG, while predicted entities by BERTRL with fewer hops spread more widely on a KG.
5 Proposed Approach
Figure 2: An overview of ReDistLP.
5.1 Model details
In this section, we introduce ReDistLP, a re-ranking model for predicting missing links. Figure 2 illustrates an overview of our model. ReDistLP is a cascaded model that first scores candidate entities using an initial retriever, and then re-ranks them with a re-ranker. As discussed in Section 3, we aim to select the initial retriever and re-ranker models such that their prediction sets ($S_{m_{1}}$ and $S_{m_{2}}$) satisfy the condition ($S_{m_{1}}\not\subset S_{m_{2}}$), according to equation (7). This non-subset relationship between the prediction sets can enhance the effectiveness of the re-ranking process. As shown in the analysis in Section 4, when trained with longer hop rules, the BERTRL tends to predict entities concentrated within narrower regions of the KG. To satisfy the condition in equation (7) and maximize re-ranking effectiveness, we employ a BERTRL trained on shorter hop rules as the initial retriever, and a BERTRL trained on longer hop rules as the re-ranker Therefore, in our final ReDistLP, we use BERTRL trained with 2-hop as the initial retriever and BERTRL trained with 3-hop as the re-ranker.
We make two variants of datasets for BERTRL in training and inference. The first is the same as the original one. This variants dataset excludes the triples where the head and tail entities do not connect within $n$-hop. The second variant does not exclude such triple data. Table 3 shows examples of training data. The first dataset only contains No.1 and 2 in this case, while the second dataset contains all cases. This enables the model to score entities that do not connect to the other entity in a triple. We call BERTRL tested on the second dataset as BERTRL v2. We use BERTRL v2 as the initial retriever for our model.
Table 3: Examples of training data from NELL-995
No.1 Question: country : republic country also known as what ? Is the correct answer country : vietnam ? Context: country : republic country located in geopolitical location state or province : states; state or province : states state located in geopolitical location country : vietnam;
No.2 Question: country : republic country also known as what ? Is the correct answer country : vietnam ? Context: country : republic country located in geopolitical location state or province : states; country : vietnam country located in geopolitical location state or province : states;
No.3 Question: country : republic country also known as what ? Is the correct answer hospital : hialeah ? Context: None
No.4 Question: country : republic country also known as what ? Is the correct answer city : beverly hills ? Context: None
5.2 Rank Threshold Optimization
Determining the optimal cutoff point for the initial retriever’s results directly affects the accuracy of the re-ranking model. Ideally, the cutoff would be defined as the position of the answer entity. Actually, in practice, model predictions are continuous probability scores, which can be noisy, making it difficult to identify a stable threshold. We explored different threshold optimization methods to find the most suitable approach for our model.
Constant Threshold One straightforward approach is to use a constant threshold, where a fixed threshold is applied to the ranking list. We determine the optimal fixed threshold using a development dataset and then apply it to the test set.
Top-$k$ Cutoff Another common approach is to use a top-k cutoff, where the top-$k$ items from the ranked list are selected. This method ensures a fixed number of items are always selected, regardless of their absolute scores.
k-means clustering A third approach is to use k-means clustering, where the items in the ranked list are grouped into k clusters based on their similarity. Items belonging to the top-scoring cluster(s) are selected. This method groups similar items together and selects the most relevant cluster(s) based on their overall scores.
Continuous approach The discrete methods mentioned above have a major drawback: they often exclude the answer entity from the selected set. When the answer entity is not included in the initial pool, the accuracy of the re-ranking model is limited to that of the initial retriever. In contrast, a continuous approach can be employed to mitigate this issue. Instead of applying fixed cutoff points, this method passes the results with their predicted scores to the re-ranker. The re-ranker then re-ranks the results while consulting the scores of the initial retriever. This approach allows for a more flexible selection process.
  • Mean
    This method calculates the final score by simply averaging the prediction scores from the two models. Our final design uses this method.
  • Mean of Normalized Scores (MNS)
    The confidence for a specific input may differ between the two models. To account for this, we normalize each score before averaging them.
  • Reciprocal Rank (RR)
    In this method, we calculate the reciprocal rank of each entity in the prediction sets of both models. The final score is determined by the mean of these reciprocal ranks.
  • Reciprocal Rank with Score (RRS)
    In this method, we multiply the score by the reciprocal rank for each entity in both models. The final score is determined by the mean of these values.
6 Experiments
In our experiments, we evaluate the effectiveness of our proposed model on inductive knowledge graph completion datasets (Section 6.4). In the latter part of Section 6.4, we analyze the performance of our model in the ideal setting. Subsequently, we conduct an ablation study in Section 6.5.
6.1 Datasets
We use inductive subsets of three knowledge completion datasets: WN18RR (Dettmers et al., 2018), FB15k-237 (Toutanova et al., 2015), and NELL-995 (Xiong et al., 2017), which were introduced by Teru et al. (2020). WN18RR is a subset of WordNet that stores lexical relations of words, with its inverse relations removed. FB15k-237 is a subset of Freebase that stores general knowledge facts. NELL-995 is a subset of NELL that is constructed from the web automatically. Table 4 provides a detailed overview of the statistics of KG datasets.
Table 4: Number of entities, relations, and triples in training and inference graphs.
Dataset Graph $||\mathcal{E}|$$||\mathcal{R}|$$|\mathcal{T}|$
WN18RR train 2746 9 6670
train-2000 1970 9 2002
train-1000 1362 9 1001
test-inductive 922 8 1991
FB15k-237 train 1594 180 5223
train-2000 1280 180 2008
train-1000 923 180 1027
test-inductive 1093 142 2404
NELL-995 train 2564 88 10063
train-2000 1346 88 2011
train-1000 893 88 1020
test-inductive 2086 79 6621
6.2 Baselines and Implementation Details
We compare our approach with GraIL (Teru et al., 2020), BERTRL (Zha et al., 2022), KG-BERT (Yao et al., 2019), RuleN (Meilicke et al., 2018), and KRST (Su et al., 2023). GraIL is a GNN-based model that predicts over subgraph structures. KRST is a transformer-based rule-induction method that selectively incorporates reliable rules during training. We use the implementation provided publicly by the authors and the best hyper-parameter settings reported in their work.
6.3 Experiment Settings
We report the evaluation metrics Hits@1 and Mean Reciprocal Rank (MRR). Hits@1 measures the percentage of cases where the correct triple is ranked as the top 1 prediction. MRR computes the average of the reciprocal ranks of the correct triple. Consistent with the evaluation setup in Teru et al. (2020) and other prior work, we measure the rank of the correct triple among a set of 50 candidate triples, comprising the correct triple and 49 negative triples.
6.4 Results and Analysis
Actual Setting. Table 5 and Table 6 present the Hits@1 and MRR results of inductive link prediction, respectively. We adopt the BERTRL v2 with 2-hop as the initial retriever and the BERTRL with 3-hop as the re-ranker of our final model.
Table 5: Hits@1 of link prediction in the inductive setting.
WN18RR FB15k-237 NELL-995
RuleN .745 .415 .638
GRAIL .769 .390 .554
KG-BERT .436 .341 .244
BERTRL .755 .541 .715
KRST .835 .602 .649
BERTRL (1 hop) .630 .329 .466
BERTRL (2 hop) .729 .410 .639
BERTRL (3 hop) .793 .578 .731
ReDistLP (2 to 3-hop) .796 .634 .748
Table 6: MRR of link prediction in the inductive setting.
WN18RR FB15k-237 NELL-995
RuleN .780 .462 .710
GRAIL .799 .469 .675
KG-BERT .542 .500 .419
BERTRL .792 .605 .808
KRST .902 .716 .769
BERTRL (1 hop) .630 .344 .466
BERTRL (2 hop) .726 .415 .637
BERTRL (3 hop) .718 .595 .726
ReDistLP (2 to 3-hop) .808 .715 .832
Our model achieves the best performance under Hits@1 on the FB15k-237 and NELL-995 datasets. Specifically, ReDistLP achieves significant improvements in the FB15k-237 (+3.2% for Hits@1) and NELL-995 (+3.3% for Hits@1). Under the MRR metric, our model achieves the best performance on the NELL-995 dataset and comparable results on the FB15k-237 dataset.
Ideal Setting. Table 7 presents the results of inductive link prediction in the ideal setting. In this ideal setting, we determine the cutoff based on the position of the answer entity in the initial pool. Using this optimal cutoff, we adopt the top-$k$ re-ranking algorithm.
Table 7: Hits@1 and MRR results on the ideal setting
WN18RR FB15k-237 NELL-995
Hits@1
ReDistLP (1 to 2-hop) .725 .590 .745
ReDistLP (1 to 3-hop) .784 .720 .827
ReDistLP (2 to 3-hop) .817 .705 .831
MRR
ReDistLP (1 to 2-hop) .742 .677 .818
ReDistLP (1 to 3-hop) .806 .781 .891
ReDistLP (2 to 3-hop) .836 .777 .893
The Hits@1 metrics show a large difference between the actual setting and the ideal setting. Specifically, Hits@1 of ReDistLP (2 to 3-hop) in the ideal setting is +2.6%, +7.1%, +8.3% higher than in the actual setting on the WN18RR, FB15k-237, NELL-995, respectively. This indicates if we could determine a more optimal cutoff, our model’s accuracy could be improved. In Section 6.5, we present an ablation study on the selection of the ranking strategy.
ReDistLP (2 to 3-hop) shows the best results. However, when compared with the initial retriever of BERTRL (Table 5, 6), ReDistLP (1 to 3-hop) demonstrates the largest improvements. To better understand the factors contributing to these differences in performance, we investigate the difference between their prediction sets. We calculate the following equation
$r=\frac{|S_{G}\setminus S_{R}|}{|S_{G}|},$ (9)
where $S_{G}$ and $S_{R}$ represent the prediction sets of the initial retriever and the re-ranker, respectively. This equation, which corresponds to equation (8), normalizes the difference in cardinality between the two prediction sets. A higher value of $r$ indicates a larger difference between the two sets. Table 8 shows the mean of $r$ calculated over the test set.
Table 8: Analysis of the difference in prediction sets between the initial retriever and re-ranker.
initial retriever re-ranker WN18RR FB15k-237 NELL-995
BERTRL 1-hop BERTRL 2-hop 0.192 0.222 0.190
BERTRL 1-hop BERTRL 3-hop 0.234 0.259 0.237
BERTRL 2-hop BERTRL 3-hop 0.192 0.221 0.191
Table 8 reveals that the pair of (BERTRL 1-hop, BERTRL 3-hop) consistently yields the highest values. This suggests that improving accuracy is achieved by maximizing the difference in prediction sets between the initial retriever and re-ranker.
6.5 Ablastion study
Table 9 shows the accuracy of different rank threshold optimization algorithms in ReDistLP. The averaging method achieves the best results among them. In the NELL-995 dataset, the RRS method outperforms the Mean method.
Overall, the discrete methods, such as k-means, top-k, and fixed threshold, could not outperform the continuous methods. However, from Table 7, even discrete methods have the potential to improve accuracy. This implies that determining the optimal cutoff in the initial retriever’s results is challenging in our model, as the set must be minimized while still including the correct entity. We will address this issue in the future study.
Table 9: Ablation study of ReDistLP variants (Hits@1)
WN18RR FB15k-237 NELL-995
top-$k$ cutoff ($k=5$) .737 .620 .702
top-$k$ cutoff ($k=10$) .739 .615 .708
top-$k$ cutoff ($k=20$) .749 .617 .722
top-$k$ cutoff ($k=40$) .780 .607 .736
fixed thresholds ($\theta=0.5$) .763 .615 .731
fixed thresholds ($\theta$ determined by the dev set) .794 .627 .735
k-means (1 out of 2) .775 .600 .704
k-means (3 out of 5) .775 .620 .719
Mean .796 .634 .748
Max .761 .627 .746
Min 0.794 0.600 0.721
MNS .778 .620 .741
RR .749 .593 .723
RRS .778 .627 .751
7 Conclusion
We proposed ReDistLP, a re-ranking model designed for inductive link prediction on KGs. ReDistLP employs two variants of BERTRL, which are trained with different rules. Different rules change the model’s behavior. Specifically, the model trained with longer hop rules predicts entities within narrower regions of the KG. ReDistLP leverages this phenomenon through a re-ranking pipeline to improve link prediction performance. Our model outperforms the baselines in 2 out of 3 datasets.
Furthermore, we demonstrated that our model’s performance can be further improved if an optimal rank threshold for re-ranking could be determined. In the future, we plan to optimize the rank threshold for our model.